Download
onsite status n.
Skip this Video
Loading SlideShow in 5 Seconds..
Onsite Status PowerPoint Presentation
Download Presentation
Onsite Status

Onsite Status

81 Vues Download Presentation
Télécharger la présentation

Onsite Status

- - - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - - -
Presentation Transcript

  1. Onsite Status Itaru Nakagawa RIKEN/RBRC

  2. Schedule

  3. Optical Fiber Connection • Completed by Seiko & Yoshi

  4. Fiber Chain Test • Completed chain test for data cables. Found 5 dead trunk cables. (Seiko, Yudai, Ray) • Arranged repair work with ITD tech on next Thursday (Sept.10). • Beam Clock (BCLK) cable test is now underway. (Seiko, Ray)

  5. LV Power • Completed terminal block connections • Megapack will be installed today • Remote control will be tested next • Voltage measurement and adjusting Last week frank Present MegaPack Power Supply

  6. Beam Clock (BCLK) Fanout 50 more optical drivers to be Delivered. LV Distribution Modules • 5 + spare cards to be produced (~1 month) LV distributer

  7. OASys Maintenance Camera/power dead? Bad power connector  FIXED! Confirmed Powers are back (Yoshi, Yudai) Signal cable swapped

  8. OASys Camera Magnet Effect • Tested with permanent magnet. Less than 30 micro meters moved. (Yoshi, Yudai) • Try to get magnetic field probe to measure B field of the permanent magnet. Thermocouple Install in South • Completed Sensor assembly • Ready to run cables. (Seiko)

  9. Station-1 Status • Octant-7 Chassis Mount Done (Mike Lenz) • Pluming Work Done (Mike Lenz) • Installing Boards  Cable connections/finalize cabling (Jimmy)

  10. Grounding Chassis St-2&3 • Completed Grounding Chassis. (Yudai, Ray) • Systematic modifications to original grounding scheme. (originally somewhat random configuration. ) Additional ground cable between chassis and chamber will be tested

  11. Pedestal Measurement (South, St-2) Plane-1 Plane-2 Plane-3 Plane-4 Plane-5 Plane-6 oct1 oct2 oct3 oct4 oct5 oct6 oct7 oct8 Last week (incomplete grounding) Ref. Jan.2009

  12. Pedestal Measurement (South, St-2) After grounding completed. Ref. Dec.8,2008

  13. Pedestal Measurement (South, St-3) Replaced Cathode Read Out Card (CROC) by LANL group After grounding completed. Ref. Dec.8,2008

  14. Pedestal Measurement (South, St-1) Ref. Dec.8,2008

  15. Pedestal Measurement (North,St1) Larger noise by factor of 1.5 ~ 2 since power shutdown in Aug.20 Ref: April.28, 2009

  16. Pedestal Measurement (North,St2) Ref: April.28, 2009

  17. Pedestal Measurement (North,St3) Ref: April.28, 2009

  18. Issues • Weak cooling water pressure. No flow in top of octants. FEE operation/functions?  To be modified next Tuesday by CAD techs. • John T’s Dedication on PAD chamber until next Wednesday + Mike Lenz on vacation  1 week of no progress in pluming. • DCM error&busy MuTR.S standalone (St-1&St-3). No data taking after power shutdown. • Short of LV distribution modules  Rental from North. Arrangement necessary.

  19. To complete Infrastructure • Finish Cabling (Friday) • Run MuTR.S standalone • Optical • Tray install for trunk cable loop. • Patch bay install at platform. • Trunk cable connection. • Fiber connections. • PatchBay(rackroom) - PatchBay (Platform) - Chassis Chain test • Repair dead cables. • LV • Megapack Install (Frank) • Connection to Terminal Block (Frank)  end of this week • LV measurement & tune • Chassis • Station-1, Oct-7 Mount (Mike) • Finish grounding St-2&St-3 • Plumming (John+Mike)

  20. Welcome Naruya & Junichi

  21. Plan for next week • Make MuTR.S standalone running (Itaru) • Test St-2&3 chassis-chamber ground cable (Yoshi+Ray) • LV measurement and tune (Seiko+Ray) • ADTX-FEE cable installation to St-1 (Yoshi+Itaru) • Daily calibration and analysis (Seiko+Ray) • Pedestal analysis code development (Seiko+Ray) • Board QA (Naruya+Junichi) • OASys temperature, humidity correlation study (Naruya+Junichi)

  22. Backup

  23. 4th Week September • South OASys camera alignment • Miscabling Check South camera 24, 26 • Diagnose Camera 18 • Thermocouple install to South • Develop noise chart program • Develop Cathode cable miscabling checker • OASys temperature measurement and correlation • North arm Station-3, Octant-1 noise hunt • Norm arm noise issue • FCI connector debugger (soldering?) • Rack room • Setup MRG and cabling • Run LL1 Cable • Run DCMIF-DCM cable